The sum of the 3 consecutive terms of arithmetic progression is 27.The product of the first 2 terms is 36.What are the first 3 terms of the progression?

Answer by Anushka(4)

Let first no. =x
second no. =x+y
third no. =x+2y
given,
x+x+y+x+2y=27
3x+3y=27
x+y=9

also given,
x(x+y)=36
dividing both,
x(x+y)/x+y = 36/9
x=4
y=5
Ans. First no.=4, Second no.=9, third no.=14